• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 16:05:53 on January26,2003

    Entry number 93469

    keyword=shift summary

    5.2.5 kinematics continue
    9:45 NMR measurements until 10:21
    10:25 start runs 20559/1559/1502
    11:33 start runs 20560/1660/1503
    go to 5.3.3 kinematics
    problem to control Q1 left - reset CPU required access
    13:06 start runs 20561/1561/1504 on C-12 optics 14 mkA
    in OTR 1C12 and adjust filters on 1C12 and 1H04A
    14:33 stop the run 20563/1563/1506
    move target to the reference cell position
    5 mkA, raster ON
    14:42 start runs 20564/1564/1507 on empty reference cell, beam OTR IN
    15:03 start runs 20565/1565/1508 beam OTR OUT !
    15:23 start runs 20566/1566/1509 on reference cell with N2 p = 146 psig - 145 psig
    15:47 start runs 20567/1567/1510 - on He-3 cell